snowflake0531:

@dude @jhonyy9 there's one more that idk of, healp please

snowflake0531:

prove the left equals the right side

snowflake0531:

@Angle @Tranquility

Tranquility:

This one seems relatively straightforward. If you look at both sides, what's the first thing you notice? We have something like: a^4 - b^4 = a^2 + b^2 That means if we factor the left side, we're going to get some identity that makes (a^2 - b^2) = 1 so that we're left with the RHS Can you factor \[\csc^4 x - cot^4 x\]

Tranquility:

If you need a little push in the right direction: \[ a^2 - b^2 = (a+b)(a-b)\]

snowflake0531:

[(csc x - cot x)(cot x + cot x)]^2

Tranquility:

No. \[ a^4 - b^4 = (a^2)^2 - (b^2)^2\] Now it's in the form of \[ x^2 - y^2 = (x + y)(x-y)\] \[ x = a^2 \\y = b^2\] What you wrote would end up being \[(\csc x - \cot x)^2(\cot x + \cot x)^2\] which is not equivalent to the original expression and is thus wrong

snowflake0531:

csc^4 x−cot^4 x (csc^2 - cot^2 x)(csc^2 + cot^2 x)

Tranquility:

Now apply one of the Pythagorean identities \[\cot^2\theta + 1 = \csc^2 \theta\]

snowflake0531:

Oh thank youuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuuu

Tranquility:

Anytime!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Tranquility Now apply one of the Pythagorean identities \[\cot^2\theta + 1 = \csc^2 \theta\] \(\color{#0cbb34}{\text{End of Quote}}\) explained perfect - congrats !